If 4.9g of nitrogen gas is placed in a 2 L container at 31 degrees, what is the pressure in the container?


1
Expert's answer
2022-02-24T15:36:03-0500

In Kelvins, 31oC = 31 + 273 = 304 K. According to the ideal gas law,


"PV=nRT"


"n(N_2)=4.9\ g\times\frac{1\ mol}{28.01\ g}=0.175\ mol"


"R=0.08206\frac{L\cdot{atm}}{mol\cdot{K}}" (constant)


"P=\frac{nRT}{V}=\frac{0.175\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times304\ K}{2\ L}=2.18\ atm"


Answer: 2.18 atm